# Introducing Unbiased Depth into 2D Gaussian Splatting for High-accuracy Surface Reconstruction

### v1

@@ 4.1 Experimental Setup
Implementation Details.
Building upon the 2DGS framework, we modify its CUDA kernel to output depth convergence and depth correction maps. We make a slight change to the adaptive Gaussian densification strategy, setting it to stop at 20,000 iterations. After stopping densification, we clamp the scale of Gaussians so that they do not exceed the threshold set during densification. Once the depth difference between two adjacent ray-splat intersections exceeds a threshold (set to one-fourth of the scene radius), we exclude the depth convergence loss calculation between them. In addition, we only apply gamma correction on the DTU dataset [14]. We conduct all the experiments on a single NVIDIA RTX4090 GPU with 24GB memory.
Datasets. We evaluate our method on three datasets, including DTU [14], Tanks and Temples [17], and Mip-NeRF360 [2]. The DTU dataset is a widely used benchmark for evaluating object-level 3D reconstruction quality. Each scene in the dataset contains 49 or 64 images, but does not fully cover all viewpoints of the object. Following the previous work [13, 35], we selected 15 scenes for comparison, which include both highlights and shadows, making them more challenging for reconstruction. We use Chamfer distance as the geometry metric, which measures the mean of reconstruction completeness and accuracy.
Baselines. Following [13, 35], we compare our method with state-of-the-art neural implicit surface methods, including VolSDF [32], NeuS [27], and Neuralangelo [18]. For Gaussian splatting methods, we compare with 3DGS [15], SuGaR [11], Gaussian Surfels (G-Surfel) [6], 2DGS [13], and GOF [35].
Table 1: Quantitative results of chamfer distance (mm)$\downarrow$ on DTU dataset [14]. Our method achieves the highest geometry accuracy compared to other Gaussian splatting methods. It’s also worth noting that our method surpasses the original 2DGS on all scenes.
@@
VolSDF [32] 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86 > 12h
NeuS [27] 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84 > 12h
Neuralangelo [18] 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43 0.61 > 128h
3DGS [15] 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50 1.96 8.1m
SuGaR [11] 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33 $\sim$ 1h
G-Surfel [6] 0.66 0.93 0.54 0.41 1.06 1.14 0.85 1.29 1.53 0.79 0.82 1.58 0.45 0.66 0.53 0.88 7.9m
2DGS [13] 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80 7.5m
GOF [35] 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74 23m
@@
Table 2: Quantitative comparisons on Mip-NeRF360 dataset [2]. All scores of the baseline methods are directly taken from their papers whenever available. Our method achieves the best overall performance in outdoor scenes.
@@
3DGS [15] 24.64 0.731 0.234 30.41 0.920 0.189
2DGS [13] 24.33 0.709 0.284 30.39 0.924 0.182
GOF [35] 24.82 0.750 0.202 30.79 0.924 0.184
Ours 26.54 0.795 0.226 31.04 0.920 0.200
Table 3: Ablation study of our method on DTU dataset [14].
@@
Figure 4: Qualitative comparison with 2DGS [13] on the DTU dataset [14]. Following 2DGS, meshes are extracted by applying TSDF to depth maps. The ground truth images are shown at the leftmost, followed by the normals and extracted meshes. Our method effectively fills holes in specular highlight areas and produces more complete and accurate surface reconstructions. A full qualitative comparison on DTU is included in the supplementary.
Figure 5: Qualitative comparisons with 2DGS [13] and GOF [35] on Tank and Temples [17] (the first 2 rows) and Mip-NeRF360 [2] (the last 3 rows). Our method has fixed many artifacts caused by highlight reflections visible in the results of the other two methods. Besides, compared to GOF, a 3D Gaussian-based method, 2DGS and our method produce smoother and more continuous surfaces.
@@ 4.2 Geometry Reconstruction
We first compare the geometry quality of our method on the DTU dataset [14] with all the baseline methods. As shown in Table 1, our method achieves the highest mean accuracy compared to other Gaussian splatting methods. In particular, our method consistently surpasses the original 2DGS in all scenes. Additionally, it attains comparable geometry reconstruction quality to the SOTA neural implicit methods, while offering a training speed improvement of over 100 times. Fig. 4 further presents the qualitative comparisons with 2DGS on DTU, where the improvement of our method on the holes in glossy regions is clearly shown.
We then perform a qualitative comparison of reconstruction quality on Tanks and Temples [17] and Mip-NeRF360 [2] with 2DGS [13] and GOF [35]. As shown in Fig 5, compared to the other two methods, our approach resolves many artifacts caused by specular reflections and excels in capturing sharp edges and intricate geometric details. Even when applied to outdoor scenes with larger dynamic ranges and more complex lighting, our method still performs well, fully demonstrating its robustness.
@@ 4.3 Novel View Synthesis
Although we build our method aiming at improving geometry quality, we still evaluate the quality of novel view synthesis and compare it to 3DGS [15], 2DGS [13] and GOF [35] using the Mip-NeRF360 dataset [2]. As listed in Table 2, our method significantly outperforms the previous Gaussian Splatting methods in terms of the PSNR metric, with an especially strong performance in outdoor scenes. This is because our unbiased depth estimation enforces strong constraints on the Gaussian distribution, making it better aligned with scene geometry. As a result, our optimization enhances color fidelity in detailed areas, leading to overall superior rendering quality.
Fig. 7 shows the qualitative comparisons of novel view synthesis on Mip-NeRF360 with 2DGS and GOF. Our rendered images outperform those of 2DGS with more high-frequency details and fewer artifacts on all scenes. Compared with GOF, the SOTA 3DGS-based method, our method reaches comparable renderings in planar parts (rows 2, 3, 5, and 6 in Fig. 7), while falling behind in some local dense and tangled regions (rows 1 and 4 in Fig. 7). We believe the advantage of GOF can be attributed to the use of 3D Gaussians, where small and dense 3D primitives are generated to fit these parts, leading to more visually appealing renderings. However, our method archives better PSNR in all the scenes, as shown in Fig. 7 and Table 2. We believe the advantage comes from the more accurate geometry we reconstruct, as compared in Sec.4.2.
@@
Figure 7: Qualitative comparisons of novel view synthesis with 2DGS [13] and GOF [35] on Mip-NeRF360 [2]. Our renderings exhibit more high-frequency details and fewer artifacts than 2DGS on all scenes. While it produces comparable results with GOF for most planar parts, our method falls behind in local dense and cluttered areas (rows 1 and 4). Nonetheless, our rendered images consistently achieve the best PSNR in all scenes, which we believe comes from the more accurate geometry we reconstruct.
@@ 4.4 Ablation Study
Table 3 and Fig. 6 present the ablation study on DTU dataset [14], where we progressively integrate our techniques into 2DGS to evaluate their impact on geometry quality. As shown in Fig. 6 (b), when only the depth convergence loss is applied (Table 3 B), it only partially addresses the issue of holes in glossy areas. When both of them are applied (Table 3 C), the holes on the surface are effectively resolved; see Fig. 6 (c). The full model with gamma correction achieves the best, showing improved geometric details in shadowed areas, as illustrated in Fig. 6 (d) and Table 3 D.
@@ Appendix A Evaluation metric on geometric accuracy
As the DTU dataset [14] has ground truth surface point clouds, we use Chamfer distance to evaluate the geometric quality as previous methods [13, 35]. Chamfer distance (CD) is a commonly used distance metric in fields such as point cloud registration and 3D reconstruction, which calculates the similarity between two point sets. Specifically, its calculation consists of two parts:
@@
where $P$ and $Q$ are two point cloud sets, representing the input data point cloud and the reference point cloud, respectively. The two terms above are also termed as Accuracy and Completion [13, 35], where Accuracy indicates the proximity of points in the reconstructed model to the ground truth model, and Completion assesses the extent to which points in the ground truth model are covered by the reconstructed model.
@@ Appendix B More Analysis on Depth Correction
Table 4: Ablation results of depth correction on DTU dataset [14]. We calculated the chamfer distance(mm)$\downarrow$ for each scanned scene. It can be observed that modifying the threshold of transmittance $T_{i}$ (see equation (8) of our paper, or equation (19) in Appendix B of 2DGS [13]) has a negligible impact on the reconstruction outcomes. On the other hand, the adoption of our new depth criterion significantly improves the reconstruction quality.
@@ Appendix C Additional comparisons and results
Fig. 10 show more qualitative results of our method on outdoor scenes of Mip-NeRF360 dataset [2].

### v3

@@ 4.1 Experimental Setup
Implementation Details. Building upon the 2DGS framework, we modify its CUDA kernel to output depth convergence and depth correction maps. We make a slight change to the adaptive Gaussian densification strategy, setting it to stop at 20,000 iterations. After stopping densification, we clamp the scale of Gaussians so that they do not exceed the threshold set during densification. Once the depth difference between two adjacent ray-splat intersections exceeds a threshold (set to one-fourth of the scene radius), we exclude the depth convergence loss calculation between them. In addition, we only apply gamma correction on the DTU dataset [JDV14]. We conduct all the experiments on a single NVIDIA RTX4090 GPU with 24GBytes memory.
Datasets. We evaluate our method on three datasets, including DTU [JDV14], Tanks and Temples [KPZK17], and Mip-NeRF360 [BMV22]. The DTU dataset is a widely used benchmark for evaluating object-level 3D reconstruction quality. Each scene in the dataset contains 49 or 64 images, but does not fully cover all viewpoints of the object. Following the previous work [HYC24, YSG24], we selected 15 scenes for comparison, which include both highlights and shadows, making them more challenging for reconstruction. We use Chamfer distance as the geometry metric, which measures the mean of reconstruction completeness and accuracy.
Baselines. Following [HYC24, YSG24], we compare our method with state-of-the-art neural implicit surface methods, including VolSDF [YGKL21], NeuS [WLL21], and Neuralangelo [LME23]. For Gaussian splatting methods, we compare with 3DGS [KKLD23], SuGaR [GL23], Gaussian Surfels (G-Surfel) [DXX24], 2DGS [HYC24], and GOF [YSG24].
[…]
@@
VolSDF [YGKL21] 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86 > 12h
NeuS [WLL21] 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84 > 12h
Neuralangelo [LME23] 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43 0.61 > 128h
3DGS [KKLD23] 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50 1.96 8.1m
SuGaR [GL23] 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33 $\sim$ 1h
G-Surfel [DXX24] 0.66 0.93 0.54 0.41 1.06 1.14 0.85 1.29 1.53 0.79 0.82 1.58 0.45 0.66 0.53 0.88 7.9m
2DGS [HYC24] 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80 7.5m
GOF [YSG24] 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74 23m
@@
Table 2: Quantitative comparisons on Mip-NeRF360 [BMV22]. All scores of the baseline methods are directly taken from their papers (whenever available).
@@
3DGS [KKLD23] 24.64 0.731 0.234 30.41 0.920 0.189
2DGS [HYC24] 24.33 0.709 0.284 30.39 0.924 0.182
GOF [YSG24] 24.82 0.750 0.202 30.79 0.924 0.184
Ours 24.60 0.728 0.255 30.00 0.918 0.185
Table 3: Ablation study of our method on DTU dataset [JDV14].
@@
Figure 4: Qualitative comparison with 2DGS [HYC24] on the DTU dataset [JDV14]. Following 2DGS, meshes are extracted by applying TSDF to depth maps. The ground truth images are shown at the leftmost, followed by the normals and extracted meshes. Our method effectively fills holes in specular highlight areas and produces more complete and accurate surface reconstructions. A full qualitative comparison on DTU is included in the supplementary.
Figure 5: Qualitative comparisons with 2DGS [HYC24] and GOF [YSG24] on Tanks & Temples [KPZK17] (the first 2 rows) and Mip-NeRF360 [BMV22] (the last 3 rows). Our method has fixed many artifacts caused by highlight reflections visible in the results of the other two methods. Besides, compared to GOF, a 3DGS-based method, 2DGS and our method produce smoother and more continuous surfaces.
@@ 4.2 Geometry Reconstruction
We first compare the geometry quality of our method on the DTU dataset [JDV14] with all the baseline methods. As shown in Table 1, our method achieves the highest mean accuracy compared to other Gaussian splatting methods. In particular, our method consistently surpasses the original 2DGS in all scenes. Additionally, it attains comparable geometry reconstruction quality to the SOTA neural implicit methods, while offering a training speed improvement of over 100 times. Fig. 4 further presents the qualitative comparisons with 2DGS on DTU, where the improvement of our method on the holes in glossy regions is clearly shown.
We then perform a qualitative comparison of reconstruction quality on Tanks and Temples [KPZK17] and Mip-NeRF360 [BMV22] with 2DGS [HYC24] and GOF [YSG24]. As shown in Fig 5, compared to the other two methods, our approach resolves many artifacts caused by specular reflections and excels in capturing sharp edges and intricate geometric details. Even when applied to outdoor scenes with larger dynamic ranges and more complex lighting, our method still performs well, fully demonstrating its robustness.
@@ 4.3 Novel View Synthesis
Although we build our method aiming at improving geometry quality, we still evaluate the quality of novel view synthesis and compare it to 3DGS [KKLD23], 2DGS [HYC24], and GOF [YSG24] using the Mip-NeRF360 dataset [BMV22]. As listed in Table 2, our method shows slightly better performance than the original 2DGS. The convergence constraint compels Gaussian kernels to distribute along object surfaces, which benefits geometric reconstruction but potentially compromises rendering quality. As demonstrated in section 3.2, the model exploits geometric misalignment to reconstruct object specular highlights. When such misalignment is eliminated, the smooth spherical harmonics (SH) representation struggles to represent high-frequency color variations.
Fig. 7 shows the qualitative comparisons of novel view synthesis on Mip-NeRF360 with 2DGS and GOF. Compared with GOF, our method reaches comparable renderings in planar parts (rows 2, 3, and 5 in Fig. 7), while falling behind in some local dense and tangled regions (rows 1 and 4 in Fig. 7). We believe the advantage of GOF can be attributed to the use of 3D Gaussians, where small and dense 3D primitives are generated to fit these parts, leading to more visually appealing renderings.
@@
Figure 7: Qualitative comparisons of novel view synthesis with 2DGS [HYC24] and GOF [YSG24] on Mip-NeRF360 [BMV22]. Our renderings exhibit more high-frequency details and fewer artifacts than 2DGS for most planar parts, while falling behind in local dense and cluttered areas (rows 1 and 4).
@@ 4.4 Ablation Study
Table 3 and Fig. 6 present the ablation study on DTU dataset [JDV14], where we progressively integrate our techniques into 2DGS to evaluate their impact on geometry quality. As shown in Fig. 6 (b), when only the depth convergence loss is applied (Table 3 B), it only partially addresses the issue of holes in glossy areas. When both of them are applied (Table 3 C), the holes on the surface are effectively resolved; see Fig. 6 (c). The full model with gamma correction achieves the best, showing improved geometric details in shadowed areas, as illustrated in Fig. 6 (d) and Table 3 D.
@@ 5 Conclusion
Acknowledgments
This work was supported in parts by NSFC (U21B2023), ICFCRT (W2441020), GD Basic and Applied Basic Research Foundation (2023B1515120026), DEGP Innovation Team (2022KCXTD025), SZU Teaching Reform Key Program (JG2024018), and Scientific Development Funds from Shenzhen University.
@@ Appendix A Evaluation metric on geometric accuracy
As the DTU dataset [JDV14] has ground truth surface point clouds, we use Chamfer distance to evaluate the geometric quality as previous methods [HYC24, YSG24]. Chamfer distance (CD) is a commonly used distance metric in fields such as point cloud registration and 3D reconstruction, which calculates the similarity between two point sets. Specifically, its calculation consists of two parts:
@@
where $P$ and $Q$ are two point cloud sets, representing the input data point cloud and the reference point cloud, respectively. The two terms above are also termed as Accuracy and Completion [HYC24, YSG24], where Accuracy indicates the proximity of points in the reconstructed model to the ground truth model, and Completion assesses the extent to which points in the ground truth model are covered by the reconstructed model.
@@ Appendix B More Analysis on Depth Correction
[…]
@@ Appendix C Additional comparisons and results
Fig. 10 show more qualitative results of our method on outdoor scenes of Mip-NeRF360 dataset [BMV22].